OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4 ANSWERS 1 −1 −5 −25

OpenStudy (marissalovescats):

Use the distributive property and then combine like terms to get x by itself

hero (hero):

3(x + 1) = -2(x - 1 + 2) 3(x + 1) = -2(x + 1) 3(x + 1) + 2(x + 1) = 0 5(x + 1) = 0 (x + 1) = 0/5 Finish the rest.

OpenStudy (anonymous):

thanks hero and marissalovescats

OpenStudy (marissalovescats):

Yw :')

hero (hero):

So which is correct @dropmag1234 ?

OpenStudy (anonymous):

1? @Hero

OpenStudy (anonymous):

is it 1 @Hero

hero (hero):

You're guessing. I gave you enough steps to figure it out.